OpenStudy (montverde):

A heated piece of metal cools according to the function c(x) = (.5)^(x − 11), where x is measured in hours. A device is added that aids in cooling according to the function h(x) = −x − 3. What will be the temperature of the metal after 5 hours?

OpenStudy (montverde):

@jim_thompson5910 @ganeshie8 @mathmale

jimthompson5910 (jim_thompson5910):

I'm not that familiar with physics so I don't know how they want you to combine c(x) and h(x). Do they want you to add the functions?

OpenStudy (montverde):

this is algebra 2 hmm

OpenStudy (montverde):

do you think we should add them togehter?

jimthompson5910 (jim_thompson5910):

maybe, that's just a guess though

OpenStudy (montverde):

−8° Celsius 26° Celsius 32° Celsius 56° Celsius

OpenStudy (montverde):

those are my option

jimthompson5910 (jim_thompson5910):

you might be better off asking this in physics

OpenStudy (montverde):

i dont know how to do what he explaind

jimthompson5910 (jim_thompson5910):

I guess we could do this c(x) = 0.5^(x-11) c(5) = 0.5^(5-11) c(5) = 64 h(x) = -x-3 h(5) = -5-3 h(5) = -8 p(x) = c(x) + h(x) p(5) = c(5) + h(5) p(5) = (64) + (-8) p(5) = 56 but again this is just a guess that you add the two functions

OpenStudy (montverde):

oh i got it 56

OpenStudy (montverde):

yes got it right